OpenStudy (anonymous):

can some1 help me to find the domain of the following function f(x)=log(logx)

OpenStudy (anonymous):

show the work please

OpenStudy (anonymous):

Yeah I can Now \(logx > 0\)

OpenStudy (anonymous):

If \(logx > 0 \) \(\text{Then}\) \(x > e^0\) \[x>1\]

OpenStudy (anonymous):

Don't get it! how do you get at x > e0?

OpenStudy (anonymous):

\(log_ex = 2\)\[\text{Then}\]\[x = e^2\]

OpenStudy (anonymous):

He raises both sides to e. But since it's log which is base 10 it should be: \[logx>0\] \[10^{logx}>10^0\] Log cancels with 10: \[x>1\] Since x^0 always = 1.

OpenStudy (anonymous):

The domain is the allowed x values. log(x) may not have x negative. (looking at the outermost log) Thus log(x)>0. Thus x>1

OpenStudy (anonymous):

Blah, dumb line breaks.

OpenStudy (anonymous):

\[Let\ f(a) = log(a)\]\[\implies Domain_f : a > 0\] \[\large \therefore f(log(x)) = log(log(x))\]\[\large\implies Domain_{f(log(x))}: log(x) > 0\]\[\large \implies x > 10^0 \implies x > 1\]

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

Why I cannot do the following? logx>0 puting this expression into a exponential one \[10^{0} > x\] so x<1 I know x>1 is correct. But tell me how to figure out the other way is not correct?

OpenStudy (anonymous):

You flipped the sign. For inequalities the only time the sign flips is if you divide by negative numbers.

OpenStudy (anonymous):

So when you moved the x to the right the sign should have followed it as such: \[10^0<x\]

OpenStudy (anonymous):

ahhh got it ;)